Session: Feb 2005 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (mauro at Wed Mar 9 21:20:57 2005) Success :
All scan recorded. Clear sky.

Torun (K. Borkowski at Thu Mar 10 06:47:47 2005) Minor failures :
BBC3 unlocked all the time (as in the previous experiment). The very last scan (about 20 min.) lost because data did not fit into the disk-pack loaded while the other pack was full. No other known problems.

Jodrell_Bank (Alastair Gunn at Thu Mar 10 22:26:06 2005) Success :
The Lovell Telescope was used. The schedule was altered to limit the number of phase calibrator source changes; some scans will not fringe. No known problems.

Cambridge (Alastair Gunn at Thu Mar 10 22:26:42 2005) Success :
No known problems.

Effelsberg (A. Kraus at Fri Mar 11 12:55:27 2005) Success :

Westerbork (Tony Foley at Fri Mar 11 13:48:21 2005) Minor failures :
Wb used in phased array mode. No known problems except RT4 out (oscillating receiver)

Onsala (Michael Lindqvist at Thu Mar 17 11:30:32 2005) Success :
No known problems.

Noto (P. Cassaro at Mon Mar 21 10:19:03 2005) Minor failures : POIN
Some pointing problems from 12:50 to 13:30 UT due to high elevation.